The length of a rectangle is 19 feet longer than the width. If the perimeter is 46 feet, which equation could be used to find the dimensions of the rectangle?


Answer by mananth(16946)

width =x
length = x+19
Perimeter = 2(L+W)
Perimeter = 2(x+x+19)


2(2x+19)=46
4x+38=46
4x=46-38
4x=8
/4
x=2
width = 2 feet
length = 21 feet